OpenStudy (anonymous):

Find the remainder when x3 - 5x2 + 10x - 15 is divided by each binomial. a. x - 1 b. x + 3 c. x - 5 d. x +2

OpenStudy (anonymous):

here let f(x)=x3 - 5x2 + 10x - 15 now for a we need to find f(1) f(1)=1-5+10-15=-9(remainder when the above expression is divided by x-1) for the rest find f(-3),f(5) and f(-2) respectively

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You get the same answer if you do synthetic division for each factor: E.g. x-1=0, x=1 or k=1 1 )1 -5 10 -15 1 -4 6 -4 6 -9 Add the 2 rows above, -9 is the remainder.

OpenStudy (anonymous):

thank you :) this makes more sense now